OpenStudy (anonymous):

an integer is two more than another integer. twice the first integer is one more than the square of the second integer. find two integers

OpenStudy (anonymous):

a=b+2, 2a=b^2+1

OpenStudy (anonymous):

oh thanks :)

OpenStudy (anonymous):

okay so actually i'm stuck, what do i do with the equations? :S

OpenStudy (anonymous):

You have to solve the system.

OpenStudy (anonymous):

a=5 b=3 works